### v3

# Large Scale Masked Autoencoding for Reducing Label Requirements on SAR Data

Abstract
Satellite-based remote sensing is instrumental in the monitoring and mitigation of the effects of anthropogenic climate change. Large scale, high resolution data derived from these sensors can be used to inform intervention and policy decision making, but the timeliness and accuracy of these interventions is limited by use of optical data, which cannot operate at night and is affected by adverse weather conditions. Synthetic Aperture Radar (SAR) offers a robust alternative to optical data, but its associated complexities limit the scope of labelled data generation for traditional deep learning. In this work, we apply a self-supervised pretraining scheme, masked autoencoding, to SAR amplitude data covering 8.7% of the Earth’s land surface area, and tune the pretrained weights on two downstream tasks crucial to monitoring climate change - vegetation cover prediction and land cover classification. We show that the use of this pretraining scheme reduces labelling requirements for the downstream tasks by more than an order of magnitude, and that this pretraining generalises geographically, with the performance gain increasing when tuned downstream on regions outside the pretraining set. Our findings significantly advance climate change mitigation by facilitating the development of task and region-specific SAR models, allowing local communities and organizations to deploy tailored solutions for rapid, accurate monitoring of climate change effects.
1 Introduction
Satellite remote sensing has fundamentally changed the way we address climate change, offering large scale, high resolution data for applications such as forest mapping Immitzer et al. [2016]Waldeland et al. [2022], wildfire monitoring Hu et al. [2021] and flood detection Tarpanelli et al. [2022]. Data from such tasks is crucial to address problems caused by climate change, but is restricted by the limitations of optical sensing. The inability of these sensors to operate at night, through cloud cover and without atmospheric interference mean that optical data is inappropriate for time sensitive tasks such as natural disaster management Tarpanelli et al. [2022]. Synthetic Aperture Radar (SAR) overcomes these limitations, providing more consistent, all-weather, day-night monitoring capabilities. These enhanced capabilities are invaluable for timely intervention in situations including extreme weather Otto et al. [2016], natural disasters Clarke et al. [2022], rapid ecological shifts Allen et al. [2015], and deforestation Moffette et al. [2021], all of which have implications for climate change.
While SAR’s robust capabilities offer a promising avenue for overcoming the challenges associated with optical sensors, it comes with its own set of complexities. The technical demands associated with processing SAR data, including aspects like coherence estimation and interferogram formation, make it challenging to apply conventional machine learning techniques. Such hurdles limit the ease of generating labeled data for supervised learning, limiting SAR’s effectiveness in automated analysis.
Self-supervised learning offers the advantage of learning directly from the input data without requiring ground truth labels. Methodologies such as those based on contrastive learning Radford et al. [2021], masked image modelling He et al. [2021] and knowledge distillation Caron et al. [2021] have achieved remarkable successes on RGB image data, yielding significant improvements in various tasks such as image classification, segmentation, and object detection Wang et al. [2022a]Yu et al. [2022], reducing dependency on labelled data. Despite these advancements, the application of self-supervised learning to SAR data remains relatively unexplored Wang et al. [2022b]. Approaches based on data fusion with other remote sensing data sources such as RGB satellite or aerial imagery exist Chen and Bruzzone [2022]Sun et al. [2023], but although these approaches can exploit the specifics of SAR data, they may not be robust to the absence of usable RGB data at night or under cloud cover. A small number of methods operating solely on SAR data exist Ren et al. [2021]Wen et al. [2021]Xu et al. [2021], but have not yet clearly shown the geographic or temporal generalisabilty often lacking in remote sensing models Safonova et al. [2023]. Applying self-supervised learning directly to the large amounts of available unlabelled SAR data would allow practitioners to circumvent the limitations posed by the absence of reliable RGB data at night or in cloudy conditions - improving accuracy and response time in areas such as disaster management and environmental monitoring. Moreover, the use of large-scale, geographically diverse data with a model large enough to accommodate it has the potential to overcome the generalisability issues that often plague remote sensing models, presenting a robust alternative solely based on SAR data.
In this work, we take a self-supervised pretraining scheme - masked autoencoding He et al. [2021] - that has been proven effective on curated RGB imagery, and apply it to polarimetric SAR data on a large set of data covering 8.7% of the Earth’s land surface. We finetune the pretrained model on two downstream tasks - vegetation cover prediction (per-image regression) and land cover classification (semantic segmentation). We show that, in all cases, pretraining improved performance on downstream tasks. We also show that models initialized with pretrained weights still outperform their randomly initialized counterparts when using substantially fewer labels. We show that the pretrained model generalised well to regions that were not seen in the pretraining set.
2 Methods
2.1 Data
Split
Our data comprises four areas of interest (AOIs) - China, the Continental United States (CONUS), Europe and South America. Of these AOIs, three comprise the pretraining set (Europe, CONUS, China). For each AOI, we divide imagery and labels into tiles of size 4480m×$\times$4480m and split the resulting tiles using geographic bands into train, validation and test sets, to avoid data leakage on contiguous tiles as much as possible (Appendix A.1). We used data from 2020 exclusively in this work to avoid the computational expense of preprocessing datasets from multiple years. Since the Earth is finite, it is feasible to pretrain a model on the entire planet, so future work should focus whether our approach also generalises temporally.
Input Data
For all tasks, we derived input data from ESA Sentinel-1 Level 1 Ground Range Detected SAR (S1GRD) amplitude data, tiled from Google Earth Engine using geetiles [^1] https://github.com/rramosp/geetiles We used seasonal averages (spring, summer, autumn, winter) in two acquisition modes and their logarithmic difference (VV, VH, VV-VH) as input, totalling 12 channels. The resolution of S1GRD imagery is approximately 10m/pixel.
Task Labels
Vegetation cover percentage labels were obtained from the Terra MODIS Vegetation Continuous Fields product (MODISVEG), available in Google Earth Engine as MOD44B.006222https://developers.google.com/earth-engine/datasets/catalog/MODIS_006_MOD44B. The resolution of the MODISVEG product is approximately 250m/pixel. We predicted the mean value of vegetation cover within each tile (percentage area covered by vegetation). Land cover classification labels were obtained from the ESA World Cover product (ESAWC), also from Google Earth Engine [^3] https://developers.google.com/earth-engine/datasets/catalog/ESA_WorldCover_v200. The resolution of the ESAWC dataset is approximately 10m/pixel, and spans 11 land cover classes. We report segmentation accuracy using mean intersection-over-union (mIoU). For both tasks, we evaluated downstream performance on one region within (Europe) and one outside (South America) the pretraining set. In all cases we trained one model from scratch and one with an encoder pretrained using masked autoencoding. We do not compare results directly to other work on the same datasets to avoid conflating performance differences due to the methods and architectures we chose when developing our model with those due to differences in input data type or SAR preprocessing differences.
2.2 Models
For pretraining we used a masked autoencoder with a ViT-B Dosovitskiy et al. [2021] encoder followed by a reconstruction decoder based on He et al. [2021]. We motivate our model selection by the observation that the original masked autoencoder behaves reasonably with minimal data augmentation He et al. [2021]. Selecting appropriate data augmentations for SAR data introduces additional complexity compared to RGB data - for example, rotation or flipping may introduce invariance to information specific to each polarisation of the instrument. A contrastive approach based on two different SAR modes Radford et al. [2021] - for example, polarimetry with two different polarisations or polarimetry and another mode such as coherence - may similarly neglect information specific to each mode. We therefore choose to omit comparison to additional methods in this short work, although future comparison remains of interest.
We applied two modifications to the model - we use 12 channels, as described in Section 2.1, and reduce the patch size from 32 to 16 - the same size in pixels as per the original implementation, but smaller relative to our input image size of $448\times 448$, therefore resulting in a longer sequence input to the transformer encoder. We motivate this change with the observation that distant pixels in remote sensing imagery are less likely to be correlated than distant pixels in a curated photograph. See Appendices A.2 and A.3 for qualitative reconstruction results and hyperparameter details.
For the MODISVEG task, we replaced the reconstruction decoder with a regression head comprising 1D convolutions, of output dimension 196, in both the sequence and hidden dimensions, followed by 3 fully connected layers of sizes {512, 256, 128}. We use ReLU Agarap [2019] activation functions between hidden layers and ELU Clevert et al. [2016] before the regression output.
For the ESAWC task, we followed SETR-PUP Zheng et al. [2021]. We increase the number of decoder layers compared to the original implementation to maintain a maximum upsampling of 2×2$\times$2 × per layer.
3 Results
ESAWC
Quantitative results for the ESAWC task are presented in Table 1, and qualitative results in Figure 1. In the fully supervised cases, both with and without pretraining and in the regions within (Europe) and outside (South America) the pretraining set, the model was skillful in classifying land cover (Best mIoU Europe: $0.533$, South America: $0.426$). For a fixed number of labels, in all cases, performance was improved by pretraining the encoder using masked autoencoding. The effect of pretraining on downstream performance increased when the downstream task inputs were from outside the pretraining set - using the South American data downstream, the model using the pretrained encoder and 10% of the labelled data (mIoU: 0.399) outperformed the randomly initialised, fully supervised model trained with 100% of available data (mIoU: 0.393). When performing the same ablation on the European data, the pretrained model using 10% of the downstream labels (mIoU: 0.508) did not outperform the randomly initialised, fully supervised model (mIoU: 0.522).
Table 1: ESA World Cover (ESAWC) segmentation accuracy reported as mIoU: European data is in the pretraining set, and South American data is unseen before the downstream task. Best results for each region in bold.
Europe South America
Scratch Pretrained Scratch Pretrained
0.1% 0.345 0.405 0.214 0.267
1% 0.438 0.475 0.290 0.360
10% 0.486 0.508 0.339 0.399
100% 0.522 0.533 0.393 0.426
Figure 1: Qualitative results for ESAWC land cover classification: Land cover classification for ESAWC on data from Europe (top row) and South America (bottom row).
MODISVEG
Quantitative results for the MODISVEG task are presented in Table 2, and correlation plots in Figure 2. In all fully supervised cases the model was skillful in predicting mean vegetation cover percentage, and pretraining using masked autoencoding improved performance in all cases. The effect of pretraining was very strong for this task - the pretrained model needed an order of magnitude less data than the randomly initialised model to achieve the same or greater performance for all data percentages in both regions. Again, the effect of pretraining increased in the region outside of the pretraining set (South America), with the pretrained model tuned using 1% of the task labels (RMSE 8.390 Veg %) outperforming the fully supervised randomly initialised model (RMSE 8.883 Veg %). For the model tuned on the European data, the pretrained model using 10% of the task labels (RMSE 3.282 Veg %) outperformed the fully supervised, randomly initialised model (RMSE 3.749 Veg %),
Table 2: Terra MODIS Vegetation (MODISVEG) prediction reported as RMSE (mean vegetation cover %): European data is in the pretraining set, and South American data is unseen before the downstream task. Best results for each region in bold.
Europe South America
Scratch Pretrained Scratch Pretrained
0.1% 7.187 5.138 14.326 10.902
1% 5.671 4.082 12.415 8.390
10% 4.171 3.282 10.116 7.256
100% 3.749 3.032 8.883 5.943
although the model tuned with 1% of the task labels did not (RMSE 4.082 Veg %). It is unclear if the improved label efficiency for regions outside the training set is due to geographic diversity or due to the encoder being trained on a larger combined set of pretraining and training tiles.
Figure 2: Correlation plots for MODISVEG prediction finetuning the pretrained model with 100% of the labelled data: Europe (left column) and South America (right column). Linear fits obtained by ordinary least squares (OLS).
4 Conclusions
Satellite remote sensing with Synthetic Aperture Radar (SAR) offers significant advantages over optical sensors, notably the ability to operate in all-weather conditions and during both day and night. These capabilities are essential for timely responses in climate change mitigation and natural disaster management. Processing and labelling this data, however, is subject to substantially more complexity. In this context, we showed that self-supervised pretraining on SAR data using masked autoencoding dramatically reduces the label requirements for effective performance in downstream tasks. The benefits of pretraining were particularly pronounced for geographic regions not seen during pretraining. By reducing label requirements and improving geographic generalisability, our work enables the application of deep learning to SAR for all-weather, day-night monitoring - significantly improving our capability to address climate change on a near-real-time basis. This enhanced monitoring frequency is crucial during extreme weather events, natural disasters, and rapid ecological changes, allowing for more timely intervention and mitigation strategies.
Acknowledgments and Disclosure of Funding
This work has been enabled by Frontier Development Lab Europe (https://fdleurope.org) a public / private partnership between the European Space Agency (ESA), Trillium Technologies, the University of Oxford and leaders in commercial AI supported by Google Cloud and Nvidia, developing open science for all Humankind. L.M-F. was supported by the European Research Council (ERC) Synergy Grant “Understanding and Modelling the Earth System with Machine Learning (USMILE)” under the Horizon 2020 research and innovation programme (Grant agreement No. 855187). M. J. A. was supported by the UKRI Centre for Doctoral Training in Application of Artificial Intelligence to the study of Environmental Risks [EP/S022961/1], and additionally by Trinity Hall, Cambridge. We are also indebted to Nicolas Longépé, Carlos López-Martínez, Fabio A. González Osorio, Samuel Bancroft, Emma Hatton, Alison Lowndes, Alistair Francis, Ioanna Bouri and the rest of reviewers during 2023 FDL-Europe sprint.
Appendix A Supplementary Material
A.1 Data Split
[…]
Figure A.1: Data split: Geographic bands for training/validation/test sets at a ratio of 60:20:20.
Table A.1: Statistics for S1GRD dataset tiles:. Regions in the pretraining set are shown in bold. Regions used in downstream tasks are shown in italics.
[…]
land surface
China 285402 5.728×$\mathbf{\times 10^{6}}$ 3.7% 1740
Conus 167403 3.360×$\mathbf{\times 10^{6}}$ 2.3% 1003
Europe 200489 4.024×$\mathbf{\times 10^{6}}$ 2.7% 1228
South America 83756 1.681×$\times 10^{6}$ 1.1% 502
Pretrain 653294 12.112×$\mathbf{\times 10^{6}}$ 8.7% 3971
Total 737050 14.793×$\times 10^{6}$ 9.8% 4473
A.2 SAR Reconstructions
Reconstructions by the masked autoencoder of SAR data masked during pretraining can be seen in Figure A.2. Note that the explicit aim of pretraining is to learn input features, not to obtain high reconstruction accuracy. The model largely predicts low-frequency features, as in He et al. [2021].
Figure A.2: Masked SAR Reconstruction: Masked autoencoder-based reconstruction of SAR amplitude imagery from the validation set. Within each row, we show the masked image (left), reconstruction (centre) and original image (right). A masking ratio of 0.75 was applied to patches of size $16\times 16$ on images of size $448\times 448$.
A.3 Training Details
Hyperparameter details for MAE-based pretraining and task finetuning can be seen in Table A.2. The patch size was halved relative to the size of the image compare to He et al. [2021], based on the intuition that distant pixels in remote sensing imagery are less likely to be correlated than distant pixels in curated photographs. This intuition appeared to be evidenced by linear probe performance on the validation data for vegetation prediction, although we do not report the results from the linear probe as it was not used on all tasks or regions. Beyond this, the probe was not used to make further design decisions. Increasing or decreasing the learning rate by an order of magnitude did not improve convergence on the validation data. No significant hyperparameter tuning was undertaken beyond these two decisions, as the computational expense of performing an equal level of tuning for all tasks and regions was too high.
Table A.2: Hyperparameter details for MAE pretraining and downstream training on MODISVEG and ESAWC.
MAE Pretraining MODISVEG ESAWC
Encoder (Params) ViT-B Dosovitskiy et al. [2021] (88.8M) ViT-B Dosovitskiy et al. [2021] (88.8M) ViT-B Dosovitskiy et al. [2021] (88.8M)
Decoder (Params) Reconstruction He et al. [2021] (5.5M) $1\times 1$ Conv + FC (Section 2.2; 19.7M) SETR-PUP Zheng et al. [2021] (3.0M)
Loss Function MSE RMSE Cross Entropy
Input Image Size $448\times 448$ $448\times 448$ $448\times 448$
Output Image Size $448\times 448$ Scalar $448\times 448$
Patch Size $16\times 16$ $16\times 16$ $16\times 16$
Masking Type Random N/A N/A
Masking Ratio 0.75 N/A N/A
Optimiser AdamW AdamW AdamW
Learning Rate 1.00E-04 1.00E-04 1.00E-04
No. Epochs (Pretrain) 75 N/A N/A
No. Epochs (0.1%) N/A 75 100
No. Epochs (1%) N/A 50 75
No. Epochs (10%) N/A 50 50
No. Epochs (100%) N/A 25 35
A.4 Training Histories
Training histories for MODISVEG and ESAWC can be seen in Figures A.3 and A.4 respectively. We did not tune the optimiser extensively, beyond changing the learning rate to achieve a reasonable rate of convergence. Note that the model was not evaluated before the first epoch (Epoch 0), so datapoints indicated at Epoch 0 are after one epoch of training.
Figure A.3: Training histories for Terra MODIS Vegetation prediction. Datapoints indicated at Epoch 0 are after one epoch of training. Models were not evaluated before the first epoch.
Figure A.4: Training histories for ESA World Cover prediction. Datapoints indicated at Epoch 0 are after one epoch of training. Models were not evaluated before the first epoch.
Aauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">